# Parse Trees Guided LLM Prompt Compression

### v1

@@ Abstract
Index Terms:
Large Language Models, Prompt Compression, Parse Trees, Prompt Structure, Text Pattern Analysis.
@@ I Introduction
Most LLMs adopt the transformer architecture, which results in the computational cost of LLMs being proportional to the square of input context length [8]. In addition, there is an upper token limit that can be processed by LLMs [8]. The use of long prompt for providing rich contexts to LLMs would also significantly increase the computational cost and might exceed the input token limit. To this end, prompt compression, by shortening the length of prompt, is suggested to reduce the computational cost during LLM inference and make LLMs possible for handling prompts beyond the input token limit (or indirectly increasing the limit).
Recently, some prompt compression methods [9, 10] feed the original prompt into a language model for generating the compressed prompt. These generative compression methods exploit the language understanding and generation ability of language models, but would encounter issues like hallucination [11] due to the limitations of generative language models. Another typical type of prompt compression methods [12, 13, 14, 15, 16] obtain the compressed prompt by selecting a portion of the original prompt without generating new contents, and thus alleviate the hallucination issue. These selective compression methods employ a measure to evaluate the importance of tokens in the original prompt, and preserve the important parts while outputting the compressed prompt.
@@
Extensive experiments are conducted on various datasets to present the performance for understanding, summarization, in-context learning, and math question answering prompts. The results demonstrate that PartPrompt considerably outperforms the state-of-the-art prompt compression methods over several compression ratios and metrics. And the superior performance of PartPrompt can be also observed when feeding the compressed prompt to different LLMs as the target model for inference. Besides, in-depth ablation experiments confirm the effectiveness of the key components in PartPrompt such as introducing parse trees, constructing the global tree, and adjusting the node value. Moreover, we further explore the performance of PartPrompt under the extreme long prompt case, and investigate the coherence of the compressed prompt via quantitative metrics and intuitive examples.
@@
  • To benefit future research, the code is freely available at https://github.com/LengendaryHippopotamus/PartPrompt
@@ II-B Prompt Compression
Some selective compression methods directly employ an existing pretrained language model to evaluate the importance of tokens. Selective-Context [12] obtains the compressed prompt by retaining tokens with higher information entropy, which is computed by a pretrained language model. Then, LLMLingua [13] further introduces the budget controller, iterative token-level compression, and distribution alignment for a better a information entropy estimation and a higher compression rate. Apart from them, other selective compression methods try to train a new model to evaluate the importance of tokens. [14] trains a model to learn the token compression given by the GPT4. [15] exploits reinforcement learning to train a model for deciding which tokens to remove, and [16] also uses reinforcement learning to train a model for pruning the prompt with the chain-of-thought style.
@@ II-C Text Pattern Analysis and Compression
Text data contains the latent patterns that can be used to assist test analysis and processing. For example, [29] employs backbone information to help the transformer model to learn the text encoding and representation. Some explicit text patterns, e.g., the patterns given by linguistic parse trees, can be easily understood by humans and have been applied to many machine learning tasks. For instances, [17] employs the parse tree to aid natural language processing tasks like machine translation. Both [18] and [19] employ parse trees to enhance the reasoning ability of models. Note that, the most related work to our work, [20] also utilizes the explicit text patterns, i.e, parse trees, to compress sentences.
@@ III Problem Formulation
Definition 2. Selective Compression refers to the compressed prompt $R_{\text{cp}}\triangleq r_{w_{1}}r_{w_{2}}\cdots r_{w_{l}}$ being selected from a part of the original prompt $R=r_{1}r_{2}\cdots r_{n}$, where $\{w_{1},w_{2},\cdots,w_{l}\}$ is a subset of $\{1,2,\cdots,n\}$ satisfying $w_{1}<w_{2}<\cdots<w_{l}$, which indicates that the selective compression does not generate new tokens and preserves the order. Accordingly, the compression ratio between the compressed prompt and original prompt can be defined by
@@
Definition 3. Selective Prompt Compression Problem can be formulated as: for a given prompt $R$ and compression ratio $\tau$ where $10<\tau<$ 1, find the compressed prompt $R_{\text{cp}}$, selecting from the original $R$, such that the overall important score or value of $R_{\text{cp}}$ is maximized without exceeding the upper token limit set by $\tau$ of $R$ during the compression process, i.e.,
@@
Definition 4. Parse Tree $T\triangleq(R,V,v_{\text{root}},f)$ consists of four terms, where $R$ is the given prompt; $V=\{v_{1},v_{2},\cdots,v_{n}\}$ is a node set for the parse tree wherein $v_{\text{root}}$ is the root node without parent node; for each node $v_{i}\in V$, function $f:V\setminus\{v_{\text{root}}\}\mapsto V$ maps each node to its parent node except for the root node. Note that, each node $v_{i}$ corresponds one-to-one to a token $r_{i}$, and hence to obey the order of tokens in the prompt, the nodes in a parse tree also retains the same order. We define $v_{w_{1}}<v_{w_{2}}$ (here $<$ indicates that it is ordered earlier in the sequence among sibling nodes) for subscript $w_{1}<w_{2}$. The node with the smallest index among child nodes is called the first child node. The length and value of a node can be defined by the corresponding token, i.e., $C(v_{i})\triangleq C(r_{i})$ and $E(v_{i})\triangleq E(r_{i})$ respectively.
Remarks: We expand the definition of the parse tree based on the foundation of traditional linguistics. If the given prompt contains only one sentence, the parse tree is exactly the traditional linguistic parse tree of the sentence. If the given prompt includes multiple sentences, the parse tree becomes the global parse tree, which integrates multiple linguistic (also referred as local) parse trees following some logic.
@@ IV The Proposed Method
This section elaborates the proposed method in detail, and the overall framework of proposed PartPrompt is illustrated in Figure 1. To be more specific, for a given prompt $R$, we first slice it into sentences $[R_{1},R_{2},\cdots,R_{m}]$ and calculate the information entropy of each token within each sentence. A parse tree is then built based on linguistic rules for each sentence so that we obtain corresponding $m$ local parse trees $[T_{1},T_{2},\cdots,T_{m}]$. Second, a global parse tree $T$ is constructed by using virtual nodes to merge these local parse trees based on the structure of prompt. Third, a node value adjustment module, including the root-ward propagation and leaf-ward propagation, is proposed to adjust the original value of each node based on human writing logic. Finally, a recursive algorithm is developed to prune the global tree, which yields the compressed tree $T_{cp}$ and accordingly the compressed prompt $R_{cp}$ for a given compress ratio $\tau$.
@@ IV-A Information Entropy Approximation
Given a prompt $R$ that contains multiple sentences, we slice it into a list of sentences $[R_{1},R_{2},\cdots R_{j},\cdots R_{m}]$. For each sentence $R_{j}$, it can further be divided into a sequence of tokens $r_{j,1}r_{j,2}\cdots r_{j,i}\cdots r_{j,n_{j}}$. The conditional probability of a token $r_{j,i}$ given its preceding token sequence can be computed by
@@
where the subscript $j$ denotes sentence index and $<j$ indicates all the tokens prior to sentence $j$, while the subscript $i$ denotes token index and $<i$ indicates all the tokens prior to the token $i$ within the sentence $j$. The corresponding information entropy [30] of a token $r_{j,i}$ then becomes
@@
As shown in Equation (4), to obtain the information entropy of a token, we can compute its conditional probability, which is exactly what the language model does. Accordingly, the information entropy of a token $r_{j,i}$ estimated by language model can be calculated by
$-\log~{}p_{\text{LM}}(~{}r_{j,i}~{}|~{}r_{j,<i},r_{<j}).$ (5)
@@ IV-B Global Parse Tree Construction
A local parse tree, based on linguistic rules, is a tree that analyzes word dependencies for a given sentence. Concretely, in a local parse tree $T_{j}$ corresponding to a sentence $R_{j}$, the word $r_{i}$, attached to the child node $v_{i}$, is considered dependent on the word of its parent node via mapping $f_{j}$ on $v_{i}$. The word of global verb in this sentence is located at the root node $v_{\text{root},j}$ without its parent dependency. We adopt Stanford NLP toolkit [31] to build all local parse trees $[T_{1},T_{2},\cdots,T_{m}]$ for all sentence $[R_{1},R_{2},\cdots,R_{m}]$.
@@ IV-C Token Alignment
The aim of token alignment is to ensure the smallest token retaining sufficient semantic integrity. To this end, we choose the tokenizer of establishing local parse trees (denoted as parse tree tokenizer) as the base, and utilize a small well-trained LLM with LLM tokenizer to compute the information entropy of the token(s) attached to actual nodes in local parse trees. The alignment between the two tokenizers achieved through a rule-based matching algorithm, yielding the aligned information entropy or value $E_{\text{Aligned}}(v_{i})$ and its length $C(v_{i})$ for each node $v_{i}$ in local parse trees.
@@ IV-D Node Value Adjustment
Section IV-B have introduced a novel global tree composing of actual nodes and virtual nodes. The hierarchical structure of virtual nodes simulates the typical writing style of sentence-paragraph-section-document. The aim of virtual nodes is to harmonize the compression requests for the segments at the global scale, where each segment refers to a subtree rooted by a virtual node. For this purpose, two pivotal criteria have to consider. First, each virtual node should reflect the value of its corresponding segment. Second, the compression request for each segment should be able to propagate to its corresponding actual nodes as well as virtual nodes.
@@
Root-ward propagation tries to update the value of each virtual node (see Lines 5 and 7) such that the value of a virtual node reflects the value of its corresponding segment. Inspired from the forward and backward propagation in neural networks, we employ a momentum-based approach to aggregate node values from leaf nodes to the root node of a segment, thereby assigning the averaged value to the corresponding virtual node. Note that, the value of actual nodes are not updated and we append the aggregated value to a list or vector $\vec{M}$, whereas the value of virtual nodes are updated recursively and we also append the aggregated value to $\vec{M}$.
Leaf-ward propagation attempts to update the value of each actual node (see Lines 17 and 18) based on its original value and the compression request. The compression request is propagated recursively from the root node to leaf nodes. For a virtual node, we employ a scalar $M$ (at very beginning $M=1$) to cache its value, and we adjust $M$ with a hyper-parameter $a_{2}$ if it is the first virtual node at a hierarchy of the global tree. For an actual node, its adjusted value is obtained by adding its original value (i.e., the aligned information entropy) and the cached $M$ with an experiential adjustment hyper-parameter $a_{1}$. Note that, $a_{1}$ ensures the adjusted value retaining a notable distinction. And $a_{2}$ targets at emphasizing the value of the first part at each hierarchy of the global tree, such as the first sentence in a paragraph and the first paragraph in a section.
@@ 11            $N\leftarrow$ Average($\vec{M}$)
12            $E_{\text{Adjusted}}(v_{i})\leftarrow$ $N$
@@ 28                  $M\leftarrow M\cdot a_{2}$
33            $E_{\text{Adjusted}}(v_{i})\leftarrow E_{\text{Aligned}}(v_{i})+M^{a_{1}}$
@@ 37            LeafwardPropagation($v_{k}$, $M$)
41LeafwardPropagation($v_{\text{root}}$, $1$)
@@ IV-E Tree Compression Based on Node Value
After obtaining the global parse tree as well as the length and the value of each node, the selective prompt compression problem, as defined in Definition 3, can be transformed into a tree pruning problem that considers the length and the value of the nodes. Formally, it can be formulated as follows.
Definition 7. Parse Tree Pruning Problem Given a tree $T=(R,V,v_{\text{root}},f)$ for a specified prompt $R$ and a specified compression ratio $\tau$ where $10<\tau<$ 1, the task is to derive a compressed tree $T_{\text{cp}}=(R_{\text{cp}},V_{\text{cp}},v_{\text{root,cp}},f_{\text{cp}})$ for a selective compression $R_{\text{cp}}$. The compressed tree is the optimal solution of the following optimization problem
@@
The objective function (8) and constraint (9) are derived from the general selective prompt compression problem. The constraint (9), which corresponds to the constraint $C(R_{\text{cp}})\leq\tau C(R)$ in the general problem, ensures that the actual compression ratio $C(V_{\text{cp}})~{}/~{}C(V)$ does not exceed the given compression ratio $\tau$, thereby allowing for precise control over the length of the compressed prompt. The objective function (8), which corresponds to objective function $\max E(R_{\text{cp}})$ in the general problem, aims to preserve the value of the nodes as much as possible.
@@
A recursive algorithm is then developed to derive the optimal solution. The pseudocode is provided in Algorithm 2. The output of algorithm is a list, denoted by $Q$, which is defined as below: $Q=[Q_{0},Q_{1},Q_{2},\dots]$, where each $Q_{l}$ represents the optimal solution limited to length $l$. To be more specific, $Q_{l}=(Q_{l,\text{1}},Q_{l,\text{2}})=(E(V_{\text{cp},l}),V_{\text{cp},l})$, with $V_{\text{cp},l}$ being the node set of the optimal compressed tree limited to length $l$. Thus, the tree constructed by node set $Q_{\lfloor\tau C(V)\rfloor,\text{2}}$ is the optimal solution of the parse tree pruning problem in Definition 7. After that, the compressed prompt is obtained by concatenating the tokens corresponding to the aforementioned nodes in sequence. In Algorithm 2, each call of function CalculateSolution recursively derives the optimal solution for the subtree rooted at $v_{i}$ from the optimal solutions of the subtrees rooted at all child nodes of $v_{i}$. This process is divided into two steps: First merging the solutions from the child nodes using the function MergeSolution (Lines 11-13), and second incorporating $v_{i}$ into the merged solution (Lines 14-18). As both of these steps maintain the optimal state, the final result is also optimal.

### v3

@@ I Introduction
Most LLMs adopt the transformer architecture, which results in the computational cost of LLMs being proportional to the square of input context length [8]. In addition, there is an upper limit to the number of tokens that can be processed by LLMs [8]. The use of long prompt for providing rich contexts to LLMs would also significantly increase the computational cost and may exceed the input token limit. To this end, prompt compression, by shortening the length of prompt, is suggested to reduce the computational cost during LLM inference and make LLMs possible for handling prompts beyond the input token limit (or indirectly increasing the limit).
Recently, some prompt compression methods [9, 10] feed the original prompt into a language model for generating the compressed prompt. These generative compression methods exploit the language understanding and generation ability of language models, but would encounter issues like hallucination [11] due to the limitations of generative language models. Another typical type of prompt compression methods [12, 13, 14, 15, 16] obtain the compressed prompt by selecting a portion of the original prompt without generating new contents, and thus alleviate the hallucination issue. These selective compression methods typically adopt a measure to evaluate the importance of tokens in the original prompt, and preserve the important parts while outputting the compressed prompt.
@@
Extensive experiments are conducted on eight datasets to present the performance across various prompt task types (including understanding, summarization, in-context learning, and math question answering) and prompt text types (including English, non-English, non-natural language, short, and long prompts). The results demonstrate that PartPrompt considerably outperforms the state-of-the-art prompt compression methods over several compression ratios and metrics. And the superior performance of PartPrompt can be also observed when feeding the compressed prompt to different LLMs as the target model for inference. Besides, in-depth ablation experiments confirm the effectiveness of the key components in PartPrompt such as introducing parse trees, constructing the global tree, and adjusting the node value. Moreover, we further explore the performance of PartPrompt under the extreme long prompt case, and investigate the coherence of the compressed prompt via quantitative metrics and intuitive examples.
@@
  • To conduct comprehensive experiments, we newly create one, re-crawl two, and re-clean one prompt compression benchmark datasets. To benefit future research, the source code along with all datasets are freely available at https://github.com/LengendaryHippopotamus/PartPrompt.
@@ II-B Prompt Compression
Some selective compression methods directly employ an existing pretrained language model to evaluate the importance of tokens. Selective-Context [12] obtains the compressed prompt by retaining tokens with higher information entropy, which is computed by a pretrained language model. Then, LLMLingua [13] further introduces the budget controller, iterative token-level compression, and distribution alignment for a better information entropy estimation and a higher compression rate. Later, LongLLMLingua [29] additionally employs a question-aware approach. Apart from them, other selective compression methods try to train a new model to evaluate the importance of tokens. LLMLingua2[14] trains a model to learn the token compression given by the GPT4. [15] exploits reinforcement learning to train a model for deciding which tokens to remove, and [16] also uses reinforcement learning to train a model for pruning the prompt with the chain-of-thought style.
@@ II-C Text Pattern Analysis and Compression
The textual data contains latent patterns that can be adopted to assist text or language analysis and processing. For example, [30] employs backbone information to help the transformer model to learn the text encoding and representation. Some explicit text patterns, e.g., the patterns given by linguistic parse trees, can be easily understood by humans and have been applied to many machine learning tasks. For instances, [17] employs the parse tree to aid natural language processing tasks like machine translation. Both [18] and [19] employ parse trees to enhance the reasoning ability of models. Note that, the most related work to our work, [20] also utilizes the explicit text patterns, i.e, parse trees, to compress sentences.
@@ III Problem Formulation
Definition 2. Selective Compression refers to the compressed prompt $R_{\text{cp}}\triangleq r_{w_{1}}r_{w_{2}}\cdots r_{w_{l}}$ being selected from a part of the original prompt $R=r_{1}r_{2}\cdots r_{n}$, where $\{w_{1},w_{2},\cdots,w_{l}\}$ is a subset of $\{1,2,\cdots,n\}$ satisfying $w_{1}<w_{2}<\cdots<w_{l}$, which indicates that the selective compression does not generate new tokens, and preserves the original order. Accordingly, the compression ratio between the compressed prompt and original prompt can be defined by
@@
Definition 3. Selective Prompt Compression Problem can be formulated as: for a given prompt $R$ and compression ratio $\tau$ where $0<\tau<1$, find the compressed prompt $R_{\text{cp}}$, selecting from the original $R$, such that the overall important score or value of $R_{\text{cp}}$ is maximized without exceeding the upper token limit set by $\tau$ of $R$ during the compression process, i.e.,
@@
Definition 4. Parse Tree $T\triangleq(R,V,v_{\text{root}},f)$ consists of four terms, where $R$ is the given prompt; $V=\{v_{1},v_{2},\cdots,v_{n}\}$ is a node set for the parse tree wherein $v_{\text{root}}$ is the root node without parent node; for each node $v_{i}\in V$, function $f:V\setminus\{v_{\text{root}}\}\mapsto V$ maps each node to its parent node except for the root node. Note that, each node $v_{i}$ corresponds one-to-one to a token $r_{i}$, and hence to obey the order of tokens in the prompt, the nodes in a parse tree also retains the same order. We define $v_{w_{1}}<v_{w_{2}}$ (here << indicates that it is ordered earlier in the sequence among sibling nodes) for subscript $w_{1}<w_{2}$. The node with the smallest index among child nodes is called the first child node. The length and value of a node can be defined by the corresponding token, i.e., $C(v_{i})\triangleq C(r_{i})$ and $E(v_{i})\triangleq E(r_{i})$ respectively.
Remarks: We expand the definition of the parse tree based on the foundation of traditional linguistics. If the given prompt contains only one sentence, the parse tree is exactly the traditional linguistic parse tree of the sentence. If the given prompt includes multiple sentences, the parse tree then becomes the global parse tree, which integrates multiple linguistic (also referred as local) parse trees following some logic.
@@ IV The Proposed Method
This section elaborates the proposed method in detail, and the overall framework of proposed PartPrompt is illustrated in Figure 1. To be more specific, for a given prompt $R$, we first slice it into sentences $R_{1},R_{2},\cdots,R_{m}]$ and calculate the information entropy of each token within each sentence. A parse tree is then built based on linguistic rules for each sentence so that we obtain corresponding $m$ local parse trees $T_{1},T_{2},\cdots,T_{m}]$. Second, a global parse tree $T$ is constructed by using virtual nodes to merge these local parse trees based on the structure of prompt. Third, a node value adjustment module, including root-ward propagation and leaf-ward propagation, is proposed to adjust the original value of each node based on the human writing logic. Finally, a recursive algorithm is developed to prune the global tree, which yields the compressed tree $T_{cp}$ and accordingly the compressed prompt $R_{cp}$ for a given compress ratio $\tau$.
@@ IV-A Information Entropy Approximation
Given a prompt $R$ that contains multiple sentences, we slice it into a list of sentences $R_{1},R_{2},\cdots R_{j},\cdots R_{m}]$. For each sentence $R_{j}$, it can further be divided into a sequence of tokens $r_{j,1}r_{j,2}\cdots r_{j,i}\cdots r_{j,n_{j}}$. The conditional probability of a token $r_{j,i}$ given its preceding token sequence can be computed by
@@
where the subscript $j$ denotes sentence index and <j<j indicates all the tokens prior to sentence $j$, while the subscript $i$ denotes token index and <i<i indicates all the tokens prior to the token $i$ within the sentence $j$. The corresponding information entropy [31] of a token $r_{j,i}$ then becomes
@@
As shown in Equation (4), to obtain the information entropy of a token, we can compute its conditional probability, which is exactly what the language model does. Accordingly, the information entropy of a token $r_{j,i}$ estimated by a language model can be calculated by
$\log~p_{\text{LM}}(~r_{j,i}~|~r_{j,<i},r_{<j}).$ (5)
@@ IV-B Global Parse Tree Construction
A local parse tree, based on linguistic rules, is a tree that analyzes word dependencies for a given sentence. Concretely, in a local parse tree $T_{j}$ corresponding to a sentence $R_{j}$, the word $r_{i}$, attached to the child node $v_{i}$, is considered dependent on the word of its parent node via mapping $f_{j}$ on $v_{i}$. The word of global verb in this sentence is located at the root node $v_{\text{root},j}$ without its parent dependency. We adopt Stanford NLP toolkit [32] to build all local parse trees $T_{1},T_{2},\cdots,T_{m}]$ for all sentence $R_{1},R_{2},\cdots,R_{m}]$.
@@ IV-C Token Alignment
The aim of token alignment is to ensure the smallest token retaining sufficient semantic integrity. To this end, we choose the tokenizer of establishing local parse trees (denoted as parse tree tokenizer) as the base, and utilize a small well-trained LLM with LLM tokenizer to compute the information entropy of the token(s) attached to actual nodes in local parse trees. The alignment between these two tokenizers is achieved through a rule-based matching algorithm, yielding the aligned information entropy or value $E_{\text{Aligned}}(v_{i})$ and its length $C(v_{i})$ for each node $v_{i}$ in local parse trees.
@@ IV-D Node Value Adjustment
Section IV-B have introduced a novel global tree composing of actual nodes and virtual nodes. The hierarchical structure of virtual nodes simulates the typical writing style of sentence-paragraph-section-document. The purpose of virtual nodes is to harmonize the compression requests for the segments at the global scale, where each segment refers to a subtree rooted by a virtual node. For this purpose, two pivotal criteria have to consider. First, each virtual node should reflect the value of its corresponding segment. Second, the compression request for each segment should be able to propagate to its corresponding actual nodes as well as virtual nodes.
@@
Root-ward propagation tries to update the value of each virtual node (Lines 5 and 7) such that the value of a virtual node reflects the value of its corresponding segment. Inspired from the forward and backward propagation in neural networks, we employ a momentum-based approach to aggregate node values from leaf nodes to the root node of a segment, thereby assigning the averaged value to the corresponding virtual node. Note that, the value of actual nodes are not updated and we append the aggregated value to a list or vector $\vec{M}$, whereas the value of virtual nodes are updated recursively and we also append the aggregated value to $\vec{M}$.
Leaf-ward propagation attempts to update the value of each actual node (Lines 17 and 18) based on its original value and the compression request. The compression request is propagated recursively from the root node to leaf nodes. For a virtual node, we employ a scalar $M$ (at very beginning $M=1$) to cache its value, and we adjust $M$ with a hyper-parameter $a_{2}$ if it is the first virtual node at a hierarchy of the global tree. For an actual node, its adjusted value is obtained by adding its original value (i.e., the aligned information entropy) and the cached $M$ with an experiential adjustment hyper-parameter $a_{1}$. Note that, $a_{1}$ ensures the adjusted value retaining a notable distinction. And $a_{2}$ targets at emphasizing the value of the first part at each hierarchy of the global tree, such as the first sentence in a paragraph and the first paragraph in a section.
@@ 11    $N\leftarrow$ Average($\vec{M}$)
12    $E_{\text{Adjusted}}(v_{i})\leftarrow$ NN
@@ 28       $M\leftarrow M\cdot a_{2}$
33    $E_{\text{Adjusted}}(v_{i})\leftarrow E_{\text{Aligned}}(v_{i})\cdot M^{a_{1}}$
@@ 37    LeafwardPropagation($v_{k}$, $M$)
41LeafwardPropagation($v_{\text{root}}$, 11)
@@ IV-E Tree Compression Based on Node Value
After obtaining the global parse tree as well as the length and the value of each node, the selective prompt compression problem, as in Definition 3, can be then re-formulated as a tree pruning problem that considers the length and the value of the nodes. Formally, it can be formulated as follows.
Definition 7. Parse Tree Pruning Problem Given a tree $T=(R,V,v_{\text{root}},f)$ for a specified prompt $R$ and a specified compression ratio $\tau$ where $0<\tau<1$, the task is to derive a compressed tree $T_{\text{cp}}=(R_{\text{cp}},V_{\text{cp}},v_{\text{root,cp}},f_{\text{cp}})$ for a selective compression $R_{\text{cp}}$. The compressed tree is the optimal solution to the following optimization problem
@@
The objective function (8) and constraint (9) are derived from the general selective prompt compression problem. The constraint (9), which corresponds to the constraint $C(R_{\text{cp}})\leq\tau C(R)$ in the general problem, ensures that the actual compression ratio $C(V_{\text{cp}})~/~C(V)$ does not exceed the given compression ratio $\tau$, thereby allowing precise control over the length of the compressed prompt. The objective function (8), which corresponds to objective function $\max E(R_{\text{cp}})$ in the general problem, aims to preserve the value of the nodes as much as possible.
@@
A recursive algorithm is then developed to derive the optimal solution. The pseudocode is provided in Algorithm 2. The output of the algorithm is a list, denoted by $Q$, which is defined as below: $Q=[Q_{0},Q_{1},Q_{2},\dots]$, where each $Q_{l}$ represents the optimal solution limited to length $l$. To be more specific, $Q_{l}=(Q_{l,\text{1}},Q_{l,\text{2}})=(E(V_{\text{cp},l}),V_{\text{cp},l})$, with $V_{\text{cp},l}$ being the node set of the optimal compressed tree limited to length $l$. Thus, the tree constructed by node set $Q_{\lfloor\tau C(V)\rfloor,\text{2}}$ is the optimal solution of the parse tree pruning problem in Definition 7. After that, the compressed prompt is obtained by concatenating the tokens corresponding to the aforementioned nodes in sequence. In Algorithm 2, each call of function CalculateSolution recursively derives the optimal solution for the subtree rooted at $v_{i}$ from the optimal solutions of the subtrees rooted at all child nodes of $v_{i}$. This process is divided into two steps: merging the solutions from the child nodes using the function MergeSolution (Lines 11-13), and incorporating $v_{i}$ into the merged solution (Lines 14-18). Since both steps maintain the optimal state, the final result should also be optimal.